# {adjustwidth}0cm-0.2cm Higgs-Induced Gravitational Waves: the Interplay of Non-Minimal Couplings, Kination and Top Quark Mass

### v1

@@ 4 Probing the top quark mass with gravitational waves
Consequently, if the top quark mass is large and allows for the negative running of the Higgs self-coupling, the GW amplitude is going to be weaker due to the intrinsically lower kination scale. On the other hand, if the top quark mass is small and the Higgs is absolutely stable, the phase transition can happen at higher scales,  [^6] The consistency of the spectator-field approximation requires the Higgs field non to backreact on the gravitational background and change the effective Planck mass, with the two conditions $\rho_{\rm tac}\ll\rho_{\phi}$ and $\xi h_{\rm tac}^{2}<0.1\times M_{P}^{2}$ imposing an upper bound on the scale of kination. Due to the typical value of ${\lambda(h_{\rm min})\approx 10^{-3}}$, the parameter space we are considering is compatible with these bounds, see the discussion in Laverda:2023uqv ; Laverda:2024qjt . up to the maximal bound on the inflationary scale $\mathcal{H}_{\rm max}=4.7\times 10^{13}\textrm{ GeV}$ Bettoni:2021zhq set by the Planck Collaboration Planck:2018jri , leading to a stronger GW signal.
The threshold separating these two scenarios is easily found from the running in (3). The solutions for $\lambda(\mu)=0$ have a bifurcation point at $\lambda_{0}=0$ which implies a simple relation
$\bar{m}_{t}=\frac{m_{h}+216.87\textrm{ GeV}}{2}\,,$ (18)
giving us the largest top quark mass $\bar{m}_{t}$ that is still compatible with the absolute stability of the SM vacuum. The best current measurement of the Higgs mass $m_{h}=125.2\pm 0.11\textrm{ GeV}$ ParticleDataGroup:2024cfk corresponds to a critical top quark mass of $\bar{m}_{t}=171.035\textrm{ GeV}$ for its central value. In principle, below this threshold, the kination scale can be as high as $\mathcal{H}_{\rm max}$ but two observational constraints must be taken into account: reheating has to be achieved at temperatures above the BBN temperature and the contribution to total energy density by GWs cannot exceed the percentage allowed by the uncertainty on the number of effective relativistic degrees of freedom at BBN. The latter can be translated into the upper bound
@@
The effect these two constraints have on the parameter space can be seen in Figure 2. The bound on the maximal inflationary scale and on the effective number of relativistic degrees of freedom tends to exclude high breaking scales, while low-scale inflationary scenarios are also excluded from reheating arguments. These constraints are universal for any value of the top quark mass, since the Higgs self-coupling converges towards its IR value $\lambda\approx 0.1$ at low scales. Dotted lines indicate the energy density of the GW spectrum for different combinations of parameters $\mathcal{H}_{\rm kin}$ and $\nu$, while the top quark mass $m_{t}=171\textrm{ GeV}$ guarantees the absolute stability of the Higgs vacuum. It is interesting to notice that for $10\lesssim\nu\lesssim 30$ the peak frequency of the GW spectrum is almost independent of the non-minimal coupling to gravity and directly correlates a particular value of the top mass to a specific value of the kination scale $\mathcal{H}_{\rm kin}$.
Figure 3: SGWB signal from a Higgs HIPT for different breaking scales compared to sensitivity curves of future GW detectors: Laser Interferometer Space Antenna (LISA) amaroseoane2017laserinterferometerspaceantenna ; Robson:2018ifk , Big Bang Observer (BBO) Crowder:2005nr ; Corbin:2005ny , UltimateDECIGO Seto:2001qf ; Yagi:2011wg ; Kawamura:2020pcg , Einstein Telescope (ET) Punturo:2010zz ; Branchesi:2023mws , Cosmic Explorer (CE) LIGOScientific:2016wof ; Reitze:2019iox and Resonant Cavities (RC) Herman:2022fau . The non-minimal coupling parameter is set to $\nu=20$ while the top quark mass $m_{t}=171\textrm{ GeV}$ ensures the absolute stability of the Higgs vacuum.
@@ 4.1 Absolutely stable electroweak vacuum
Figure 4: Energy density $\Omega_{\rm GW,0}$ and peak frequency $f_{\rm GW,0}$ of the GW spectrum at the present day for different values of the top quark mass and the non-minimal coupling parameter, with a fixed phase-transition scale $\mathcal{H}_{\rm kin}=10^{12}\text{ GeV}$. The light-blue-shaded area corresponds to a phase-transition signal within the sensitivity range of the Resonant Cavities experiment in Herman:2022fau .
The full GW signal in the case of absolute stability $m_{t}<\bar{m}_{t}$ is displayed in Figure 3 for a subset of the parameter space with $\nu=20$. Higher scales of kination produce stronger signals, as it is clear from the fitting formula for the GW amplitude (14) and its rescaling (16). The typical frequency of the spectra from the phase transition is always within the range $10^{10}-10^{11}\textrm{ Hz}$, restricting the possible detection of the peak to ultra-high frequency detectors Gatti:2024mde ; Aggarwal:2020olq ; Herman:2022fau . On the other hand, the low-frequency part of the spectrum is dominated by the super-horizon inflationary tensor perturbations that are amplified as they reenter the horizon during the period of stiff expansion Allahverdi:2020bys . Their contribution to the spectrum is given by
@@
for modes reentering during the stiff phase of kination, with $k_{\rm pivot}=0.002\text{ Mpc}^{-1}$ the pivot scale at which $\mathcal{H}_{\rm max}$ is measured Planck:2018jri , $f_{\rm pivot}$ the frequency of the pivot scale, $f_{\rm ht}$ the frequency corresponding to the Hubble scale at the end of reheating and $n_{t}$ the spectral tilt of inflationary perturbations. It is interesting to notice that the IR stochastic inflationary background is within reach of low-frequency detectors for sufficiently high kination scales. Consequently, two detection windows can simultaneously measure the kination scale in the low frequencies and the characteristics of the spectrum at higher frequencies. In this way, the full parameter space $(\mathcal{H}_{\rm kin},m_{t},\nu)$ would be reduced to $(m_{t},\nu)$ with a precise relation between $m_{t}$ and $\nu$. However, breaking this last degeneracy is a more complicated task that requires either the detection of the peak of the phase-transition signal or the detection of the knees in the spectra indicating the size of the Hubble scale at reheating. Indeed, while the peak frequency is mostly dependent on the top mass, the non-minimal coupling parameter $\nu$ is the main actor in the duration of the heating stage, see Laverda:2024qjt . The combination of a GW detection with constraints on the reheating temperature would, therefore, be able to characterise the Higgs effective potential at high scales.
@@ 4.2 Unstable electroweak vacuum
Considering now the case of $m_{t}>\bar{m}_{t}$, enforcing the stability condition in (5) sets an upper bound on the kination scale Laverda:2024qjt and on the allowed top-mass range, see Figure 4. The instability region is determined by considering only the classical fluctuations of the Higgs field without taking into account the quantum tunnelling between vacua, which could further destabilise the low-scale vacuum. A quick estimate based on the Hawking-Moss and Coleman-de Luccia instanton solutions led to the conclusion that the electroweak vacuum lifetime is primarily affected by the classical overtaking of the barrier in the effective potential Laverda:2024qjt . The task of computing a more realistic tunnelling rate in a time-dependent stochastic system with out-of-equilibrium fluctuations is an interesting but complicated problem by itself that goes beyond the scope of the present work Steingasser:2023gde ; Steingasser:2024ikl ; Garbrecht:2024end .
@@
As can be seen in Figure 4, the GW signal in the case of large top-quark masses is quenched by the small gravitational effective mass. For a fiducial value of $\mathcal{H}_{\rm kin}=10^{12}\textrm{ GeV}$ the HIPT spectrum is typically below $\Omega_{\rm GW,0}\sim 10^{-9}$ and of comparable magnitude to the SGWB of amplified inflationary perturbations. Once more, it is possible to observe the tight relation between the top quark mass and the frequency of the GW signal, as noted also in Figure 2. This feature originates from the dependence of the peak frequency on the rescattering effects taking place in the non-linear phase. Larger values of $\lambda(\mu)$ lead to a quick cascade to smaller scales that produces a strong GW signal due to the large fluctuations. In this sense, a direct effect of the running of $\lambda$ and the kination scale can leave an imprint on the GW signal.
@@
Figure 5: GW signal from the non-minimally coupled Higgs for different heating efficiencies of an independent sector $\Theta_{\text{ht}}^{\chi}$ compared to sensitivity curves of proposed GW detectors. The model parameters have been set to $\nu=20$ and $m_{t}=171\textrm{ GeV}$ for a kination scale of $\mathcal{H}_{\rm kin}=10^{12}\textrm{ GeV}$.
@@
Because of the stability bounds enforced on $\mathcal{H}_{\rm kin}$, experiments at ultra-high frequencies would be able to primarily detect the blue-tilted inflationary spectrum, scaling proportionally to $f$ instead of the typical $f^{3}$-scaling of IR modes excited by causal sources Caprini:2009fx ; Cai:2019cdl . In Figure 4, the light-blue-shaded parameter space indicates that the phase-transition signal is visible by the resonant cavity experiment in Herman:2022fau . The non-shaded area is allowed by stability constraints but it is characterised by a weaker signal generally obscured by the amplified inflationary perturbations. A precise evaluation of this situation requires a numerical analysis taking into account the horizon-reentering tensor perturbations during inflation, which act as a source term for the GW equation. This goes beyond the numerical simulations performed in Bettoni:2024ixe , where a large separation of scales between the kination scale and the typical excited momenta is present. Such condition fails typically for very inefficient tachyonic production at low energy-scales or small non-minimal coupling parameters.
4.3 Independent heating mechanism
@@
Although the tachyonic particle production in a HIPT scenario is a very efficient way to achieve a reheated Universe in non-oscillatory models of inflation, it is interesting to consider the effects associated with an independent heating mechanism on the GW signal. This is particularly important for ensuring a successful reheating at low inflationary scales, $\mathcal{H}_{\rm kin}\lesssim 10^{5.5}\text{ GeV}$, where the radiation temperature of the Higgs sector alone is generically below $T_{\rm BBN}\simeq 5\text{ MeV}$. The details of the heating mechanism are not needed, as long as there exists a radiation-like sector whose energy density is greater than the Higgs’. In so doing, the heating stage is entirely described by the new sector’s heating efficiency that we label with $\chi$, namely $\Theta_{\text{ht}}^{\chi}\equiv\rho_{\chi}(a_{\rm rad})/\rho_{\phi}(a_{\rm rad})$. This quantity is the only additional free parameter we need to consider in our analysis. Since $\Theta_{\text{ht}}^{\chi}\geq\Theta_{\text{ht}}^{h}$ by construction, the heating temperature is higher than the Higgs-only scenario,
@@
$\Omega_{\rm GW,0}(\mathcal{H}_{\rm kin},\nu,m_{t},\Theta_{\text{ht}}^{\chi})=1.67\times 10^{-5}h^{-2}\left(\frac{100}{g_{*}^{\rm ht}}\right)^{1/3}\left(\frac{\Theta_{\text{ht}}^{\rm h}}{\Theta_{\text{ht}}^{\chi}}\right)\times\bar{\Omega}_{\rm GW}\,.$ (23)
Notice that we recover the Higgs-only result if we assume the Higgs to be the only reheating sector $\Theta_{\rm ht}^{h}=\Theta_{\rm ht}^{\chi}$. In Figure 5, the combined effects of an independent heating sector are visible in the relative positioning of the GW spectra. All model parameters have been fixed except from $\Theta_{\rm ht}^{\chi}$, which can take on values much larger than $\Theta_{\rm ht}^{h}$. For highly efficient heating, the spectrum is peaked at lower frequencies but at the same time its amplitude is reduced, since the Higgs is energetically subdominant with respect to the external heating sector. The introduction of the $\chi$-sector breaks the scaling relations that characterise the Higgs-only scenario by introducing a new independent energy scale. Since the duration of kination can be shortened, the amplification of the inflationary tensor perturbations is reduced and the phase-transition signal becomes more pronounced over the inflationary background the more efficient the heating sector is. The potential detection of a peaked GW signal at frequencies below the typical range in Figure 3 could be a smoking-gun clue for the existence of an external heating sector beyond the SM.

### v2

@@ 5 Probing the top-quark mass with gravitational waves
Consequently, if the top-quark mass is large and allows for the negative running of the Higgs self-coupling, the GW amplitude is going to be weaker due to the intrinsically lower kination scale. On the other hand, if the top-quark mass is small and the Higgs is absolutely stable, the phase transition can happen at higher energy scales,  [^6] Due to the typical value of $\lambda(h_{\rm min})\approx 10^{-3}$}, the parameter space we are considering is generally compatible with the condition $\Delta_{1}\ll 1$ controlling the actual onset of the HIPT, see the discussion in Laverda:2023uqv ; Laverda:2024qjt . On the other hand, the condition $\Delta_{2}\ll 1$ is violated in limited regions of the parameter space where ${\cal H}_{\rm kin}\gtrsim 10^{12}$ and $\nu\gtrsim 10-15$, indicating the breaking of the effective field theory approach. For further details, we refer the reader to Appendix C. up to the maximal bound on the inflationary scale $\mathcal{H_{\rm inf}}\leq{\cal H}_{\rm max}=4.7\times 10^{13}\textrm{ GeV}$ Bettoni:2021zhq set by the Planck Collaboration Planck:2018jri , leading to a stronger GW signal.
Figure 3: SGWB signal from a Higgs HIPT for different breaking scales compared to sensitivity curves of proposed GW detectors: Laser Interferometer Space Antenna (LISA) amaroseoane2017laserinterferometerspaceantenna ; Robson:2018ifk , Big Bang Observer (BBO) Crowder:2005nr ; Corbin:2005ny , UltimateDECIGO Seto:2001qf ; Yagi:2011wg ; Kawamura:2020pcg , Einstein Telescope (ET) Punturo:2010zz ; Branchesi:2023mws , and Cosmic Explorer (CE) LIGOScientific:2016wof ; Reitze:2019iox . The non-minimal coupling parameter is set to $\nu=20$ while the top-quark mass $m_{t}=170.8\textrm{ GeV}$ ensures the absolute stability of the Higgs vacuum.
The threshold separating these two scenarios is easily found from the running in (10). The solutions for $\lambda(\mu)=0$ have a bifurcation point at
$\bar{m}_{t}=170.886\textrm{ GeV},\qquad\mu_{\rm inst}=4.13\times 10^{17}\textrm{ GeV}\,$, (25)
giving us the largest top-quark mass $\bar{m}_{t}$ that is still compatible with the absolute stability of the SM vacuum. Below this threshold, the kination scale can be a priori as high as $\mathcal{H}_{\rm max}$ but several constraints must be taken into account: first, the self-consistency of the effective field theory approach enforces us to establish an upper cut-off $M_{P}/\xi$ on the amplitude of the Higgs fluctuations Barbon:2009ya ; Burgess:2009ea ; Hertzberg:2010dc ; Burgess:2010zq ; Bezrukov:2010jz ; Bezrukov:2012hx ; Ren:2014sya ; Bezrukov:2014ipa ; Fumagalli:2017cdo ; Antoniadis:2021axu ; Mikura:2021clt ; Ito:2021ssc ; Escriva:2016cwl ; Karananas:2022byw ; second, the heating of the Universe and the associated onset of radiation domination must be achieved at temperatures above the BBN temperature; third, the contribution of GWs to the total energy density cannot exceed the percentage allowed by the uncertainty on the effective number of relativistic degrees of freedom at BBN Allahverdi:2020bys ; Caprini:2018mtu ; Bettoni:2021zhq ,
@@
The combined effect of these consistency and viability constraints can be seen in Figure 2. High symmetry-breaking scales are excluded by the CMB bound and the self-consistency of the scenario, where the light-blue-shaded area excludes the parameter space leading to amplitudes $\langle h^{2}_{\rm br}\rangle>10^{-2}\;M_{P}^{2}/\xi^{2}$, while low-scale inflationary models are excluded from heating arguments. Notice that the heating constraint is universal for any value of the top-quark mass, since the Higgs self-coupling converges towards its IR value $\lambda\approx 0.1$ at low-energy scales. Dotted lines indicate the energy density of the GW spectrum for different combinations of the parameters $\mathcal{H}_{\rm kin}$ and $\nu$, while the top-quark mass $m_{t}=170.8\textrm{ GeV}$ guarantees the absolute stability of the SM vacuum. Interestingly, for $10\lesssim\nu\lesssim 30$ the peak frequency of the GW spectrum is almost independent of the non-minimal coupling to gravity and directly correlates a particular value of the top-quark mass to a specific value of the kination scale $\mathcal{H}_{\rm kin}$.
@@ 5.1 Absolutely stable electroweak vacuum
Figure 4: Energy density $h^{2}\Omega_{\rm GW,0}$ and peak frequency $f_{\rm GW,0}$ of the GW spectrum at the present day for different values of the top-quark mass and the non-minimal coupling parameter, with a fixed phase-transition scale $\mathcal{H}_{\rm kin}=10^{12}\text{ GeV}$. Red regions are forbidden because classically unstable, while light-blue shaded areas do not satisfy the 1% upper bound on the gravitational theory cut-off scale $\langle h^{2}\rangle<10^{-2}M_{P}^{2}/\xi^{2}$.
The full GW signal in the case of absolute stability $m_{t}<\bar{m}_{t}$ is displayed in Figure 3 for a subset of the parameter space with $\nu=20$. Higher scales of kination produce stronger signals, as it is clear from the fitting formula for the GW amplitude (21) and its rescaling (23). The typical frequency of the spectra from the phase transition is always within the range $10^{10}-10^{11}\textrm{ Hz}$, restricting the possible detection of the background to ultra-high frequency detectors. [^7] For a comprehensive analysis on the prospects of high-frequency GW detectors, see e.g. Gatti:2024mde ; Aggarwal:2020olq ; Aggarwal:2025noe , and TitoDAgnolo:2024uku for a recent evaluation of their intrinsic limitations. On the other hand, the low-frequency part of the spectrum is dominated by the super-horizon inflationary tensor perturbations that are amplified as they reenter the horizon during the period of stiff expansion Allahverdi:2020bys . Their contribution to the spectrum is given by
@@
for modes reentering during the stiff phase of kination, with $k_{\rm pivot}=0.002\text{ Mpc}^{-1}$ the pivot scale at which $\mathcal{H}_{\rm max}$ is measured Planck:2018jri , $f_{\rm pivot}$ the frequency of the pivot scale, $f_{\rm ht}$ the frequency corresponding to the Hubble scale at the end of heating and $n_{t}$ the spectral tilt of inflationary perturbations. It is interesting to notice that the IR stochastic inflationary background could be within reach of low-frequency detectors for sufficiently high kination scales, even though in this case the self-consistency of the procedure becomes potentially compromised. In this regime, two detection windows could simultaneously measure the kination scale in the low frequencies and the characteristics of the spectrum at higher frequencies. In this way, the full parameter space $(\mathcal{H}_{\rm kin},m_{t},\nu)$ would be reduced to $(m_{t},\nu)$ with a precise relation between $m_{t}$ and $\nu$. However, breaking this last degeneracy is a more complicated task that requires either the detection of the peak of the phase-transition signal or the detection of the knees in the spectra indicating the size of the Hubble scale at heating. Indeed, while the peak frequency is mostly dependent on the top-quark mass, the non-minimal coupling parameter $\nu$ is the main actor in the duration of the heating stage, see Laverda:2024qjt . The combination of a GW detection with constraints on the (re)heating temperature would, therefore, be able to characterise the Higgs effective potential at high scales.
@@ 5.2 Unstable electroweak vacuum
For top-quark masses $m_{t}>\bar{m}_{t}$, enforcing the stability condition in (12) sets an upper bound on the kination scale Laverda:2024qjt and on the allowed top-mass range, see Figure 4. The instability region is determined by considering only the classical fluctuations of the Higgs field without taking into account the quantum tunnelling between vacua, which could further destabilise the low-scale vacuum. A quick estimate based on the Hawking-Moss and Coleman-de Luccia solutions led to the conclusion that the electroweak vacuum lifetime is primarily affected by the classical overtaking of the barrier in the effective potential Laverda:2024qjt . The task of computing a more realistic tunnelling rate in a time-dependent stochastic system with out-of-equilibrium fluctuations is an interesting but complicated problem by itself that goes beyond the scope of the present work Steingasser:2023gde ; Steingasser:2024ikl ; Garbrecht:2024end .
@@
As can be seen in Figure 4, the GW signal in the case of large top-quark masses is quenched by the small gravitational effective mass. For a fiducial value of $\mathcal{H}_{\rm kin}=10^{12}\textrm{ GeV}$} the HIPT spectrum is typically below $h^{2}\Omega_{\rm GW,0}\sim 10^{-9}$ and of comparable magnitude to the SGWB of amplified inflationary perturbations. Once more, it is possible to observe the tight relation between the top-quark mass and the frequency of the GW signal, as noted also in Figure 2. This feature originates from the dependence of the peak frequency on the rescattering effects taking place in the non-linear phase. Larger values of $\lambda(\mu)$ lead to a quick cascade to smaller scales that produces a GW signal at higher frequencies. In this sense, a direct effect of the running of $\lambda$ and the kination scale can leave an imprint on the GW signal.
@@
Figure 5: GW signal from the non-minimally coupled Higgs for different heating efficiencies of an enhanced heating sector $\Theta_{\text{ht}}^{\chi}$ compared to sensitivity curves of proposed GW detectors. The model parameters have been set to $\nu=20$ and $m_{t}=170.8\textrm{ GeV}$ for a kination scale of $\mathcal{H}_{\rm kin}=10^{12}\textrm{ GeV}$.
@@
Because of the stability bounds enforced on $\mathcal{H}_{\rm kin}$, the frequency range going from kHz to GHz is dominated by the blue-tilted inflationary spectrum, scaling proportionally to $f$ instead of the typical $f^{3}$-scaling of IR modes excited by causal sources Caprini:2009fx ; Cai:2019cdl . Indeed, in Figure 4 the non-shaded area is allowed by stability constraints but it is characterised by a weaker signal generally obscured by the amplified inflationary perturbations. A precise evaluation of this situation requires a numerical analysis taking into account the horizon-reentering tensor perturbations during inflation, which act as a source term for the GW equation. This goes beyond the numerical simulations performed in Bettoni:2024ixe , where a large separation of scales between the kination scale and the typical excited momenta is present. Such condition fails typically for very inefficient tachyonic production at low energy-scales or small non-minimal coupling parameters.
5.3 Enhanced heating mechanism
@@
Although the tachyonic particle production in a HIPT scenario is a very efficient way to achieve a reheated Universe in non-oscillatory models of inflation, it is interesting to consider the effects on the GW signal associated to a different heating mechanism. This is particularly important for ensuring a successful heating at low inflationary scales, $\mathcal{H}_{\rm kin}\lesssim 10^{5.5}\text{ GeV}$, where the radiation temperature of the Higgs sector alone is generically below $T_{\rm BBN}\simeq 5\text{ MeV}$. The specific details of a fully independent heating mechanism are not needed, as long as there exists a radiation-like sector whose energy density is greater than the Higgs’. Even within our framework, two effects can lead to a larger heating efficiency than the one expected from the phase transition alone. In our starting assumptions we have neglected the additional effect that a tree-level coupling to the inflaton field might have on the Higgs fluctuations. For a generic quartic interaction $\sim g^{2}h^{2}\phi^{2}$ the non-adiabatic transition to kination can produce an initial abundance of Higgs particles Bettoni:2021qfs thus leading to overall higher Higgs energy density.  [^8] To ensure that the Higgs mass remains independent of the inflaton field in the late Universe, one must assume that, following the transition from inflation to kination, this quartic interaction either becomes negligible or is forbidden by symmetry considerations Bettoni:2021qfs ; Rubio:2017gty . On top of this effect, the non-instantaneous change in the Ricci scalar can also lead to some transient amplification of the Higgs fluctuations, a factor that is more important for small values of $\xi$.
Regardless of its fundamental nature, the heating efficiency language allows us to parametrise the macroscopic characteristics of an enhanced reheating sector, which we label with $\chi$, whose heating efficiency is simply $\Theta_{\text{ht}}^{\chi}\equiv\rho_{\chi}(a_{\rm rad})/\rho_{\phi}(a_{\rm rad})$. This quantity is the only additional free parameter we need to consider in our analysis. Since $\Theta_{\text{ht}}^{\chi}\geq\Theta_{\text{ht}}^{h}$ by construction, the (re)heating temperature for an enhanced heating is higher than the Higgs’ HIPT scenario,
@@
$\Omega_{\rm GW,0}(\mathcal{H}_{\rm kin},\nu,m_{t},\Theta_{\text{ht}}^{\chi})=1.67\times 10^{-5}h^{-2}\left(\frac{100}{g_{*}^{\rm ht}}\right)^{1/3}\left(\frac{\Theta_{\text{ht}}^{h}}{\Theta_{\text{ht}}^{\chi}}\right)\times\bar{\Omega}_{\rm GW}\,.$ (30)
Notice that we recover the Higgs HIPT result if we assume $\Theta_{\rm ht}^{h}=\Theta_{\rm ht}^{\chi}$. In Figure 5, the combined effects of an augmented heating sector are visible in the relative positioning of the GW spectra. All model parameters have been fixed except from $\Theta_{\rm ht}^{\chi}$, which can take on values much larger than $\Theta_{\rm ht}^{h}$. For highly efficient heating, the spectrum is peaked at lower frequencies but at the same time its amplitude is reduced, since the HIPT Higgs is energetically subdominant with respect to the more efficient heating scenario. The $\chi$-sector breaks the scaling relations that characterise the Higgs HIPT scenario by introducing a new independent energy scale. The potential detection of a peaked GW signal at frequencies below the typical range in Figure 3 could be a smoking-gun clue for additional new physics related to direct inflaton couplings or external heating sectors beyond the SM.